a rectangular parking lot is twice as long as it width if the perimeter of the lot is 102 find the length and width of the parking lot

Answer by macston(5194)

W=width; L=length=2W; P=perimeter=102
P=2(L+W) Substitute for L.
102=2(2W+W)
102=2(3W)
102=6W Divide each side by 6.
17=W ANSWER 1: The width is 17.
L=2W=2(17)=34 ANSWER 2: The length is 34.
CHECK:
P=2(L+W)
102=2(34+17)
102=2(51)
102=102